OpenStudy (anonymous):

PLEASE HELP. Medal! x+2 y+9 x If ---- = ----, then -- = ? 2 9 2

OpenStudy (yttrium):

Okay.

OpenStudy (yttrium):

Basically, the above equation can be rewritten as: \[\frac{ x }{ 2 } +1 = \frac{ y }{ 9 } + 1\] and as you can see, 1 will be cancelled out. Therefore, \[\frac{ x }{ 2 } = \frac{ y }{ 9 }\]

OpenStudy (anonymous):

Thank you so much! Yeah I get it now. Could you help me with more?

OpenStudy (anonymous):

@Yttrium

OpenStudy (yttrium):

Hmmm. Post it as another question. So that it will not be hassle. :)

OpenStudy (anonymous):

Ok